OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (2, 4) and has a slope of 3. y - 4 = 3(x - 2) y - 2 = 3(x - 4) y + 4 = 3(x + 2) y + 2 = 3(x + 4)

OpenStudy (anonymous):

1st one?

OpenStudy (anonymous):

point-slope form: \[\huge y-y_0=m(x-x_0) \] yes... it is the first one... you're too fast for me...

OpenStudy (anonymous):

Lol. I have another question.

OpenStudy (anonymous):

shoot...

OpenStudy (anonymous):

be right back...

OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point shown and is parallel to the line shown. The parallel line is passing through (8,5)

OpenStudy (anonymous):

y = -2x + 2 y = 2x - 14 y =1/2 x - 8 y = -1/2x - 4

OpenStudy (anonymous):

what is the slope the the parallel line?

OpenStudy (anonymous):

Idk.

OpenStudy (anonymous):

other than the point (8, 5) can you give me another point the line goes through?

OpenStudy (anonymous):

That's all it shows me.

OpenStudy (anonymous):

no graph, no picture?

OpenStudy (anonymous):

is that line falling or going up?

OpenStudy (anonymous):

Line is going up t o the left.

OpenStudy (anonymous):

so from left to right the graph is falling?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

that means it has a negative slope... the best i can do here is choose the ones with negative slope... first one and the last one...